Solve for y: y^-1/2=1/3
1. 3
2. Rad 3
3. 9
4. 27

Answer by rothauserc(4718)

Solve for y: y^-1/2 = 1/3
note that y^-1/2 = 1 / y^1/2, therefore
1 / y^1/2 = 1 / 3
cross multiply fractions
y^1/2 = 3
square both sides of =
y = 9
**************************
answer is 3. (y = 9)